OpenStudy (anonymous):

What is the t in: 3657/2720 = (1.006833333)^12t?

OpenStudy (earthcitizen):

First \[(1.006833)^{12t}=(3657)/(2720)\]

OpenStudy (anonymous):

okay.

OpenStudy (earthcitizen):

\[(1.006833)^{12t}=1.343\] taking logs of both side leaves \[12tlog(1.006833)=\log(1.343)\]

OpenStudy (earthcitizen):

divide by the log on the rhs by the lhs \[12t=\log(1.343)/\log(1.0068330=> 12t=43.31\]

OpenStudy (anonymous):

by the way, what are logs? How do I calcualte log?

OpenStudy (earthcitizen):

\[t=(43.31/12)=3.61\]

OpenStudy (earthcitizen):

logs stand for logarithm

OpenStudy (anonymous):

oh.. what is logarithm and when do I use it? (I'm so sorry for asking D:)

OpenStudy (earthcitizen):

No p! It is the power to which a number (called the base) is raised to produce a given number; e.g., the logarithm of 100 to the base 10 is 2.

OpenStudy (anonymous):

ohhh!! thank you!!

OpenStudy (earthcitizen):

no stress!